OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (anonymous):

medal and fan to someone who answers! :D

OpenStudy (anonymous):

multiply out first using the distributive law what do you get?

OpenStudy (anonymous):

im not sure could u do the steps plz? but then have me answer like right before you get the answer

OpenStudy (matlee):

can i step in

OpenStudy (ahsome):

\[3(x + 1) = -2(x - 1) - 4.\] Expand brackets \[3x+3=-2x+2-4\] Swap the places \[3x+2x=2-4+3\]\[5x=1\] \[x=\frac{1}{5}\]

OpenStudy (ahsome):

Whoops, its: \[5x=-5\]\[x=\frac{-5}{5}\]\[x=-1\]

OpenStudy (ahsome):

Hope that helps @horselover33

OpenStudy (anonymous):

so it's -1?

OpenStudy (anonymous):

@Ahsome

OpenStudy (ahsome):

Yes @horselover33 :D